OpenStudy (anonymous):

Now suppose we repeatedly select candies with replacement. Let X be the number of selections required to get the first cherry candy. What is P(X = 5)? Answer is 0.079. I want to know how to get this answer

Miracrown (miracrown):

Is this part of another problem? Something seems to be missing, we would need to know how many candies are there in total and how many of each flavor

OpenStudy (anonymous):

@Miracrown A dish contains three cherry candies, four lemon candies and five grape candies. If we randomly select two candies from the dish without replacement, what is the probability that we get one cherry candy and one grape candy? These two are together.

Miracrown (miracrown):

3 cherry candies: O O O 4 lemon candies: O O O O 5 grape candies: O O O O O How many candies are there in total? 3+4+5 = ?

OpenStudy (anonymous):

@Miracrown 12 candies. I used Binomial Distribution, but I can't get the correct answer

OpenStudy (xapproachesinfinity):

what is the prob of getting a cherry at each selection

OpenStudy (xapproachesinfinity):

since it is replacement the prob will be the same for all selection from 1 to 5

OpenStudy (xapproachesinfinity):

now you need to figure out the prob for 1 selection

OpenStudy (xapproachesinfinity):

the prob for choosing a cherry in one selection is 3/12=1/4 yes?

OpenStudy (xapproachesinfinity):

come on you need to work here!

OpenStudy (xapproachesinfinity):

how did you do it with binomial distribution?

OpenStudy (anonymous):

@xapproachesinfinity So how to get 0.079?? I know a cherry in one selection is 0.25, but I just don't know what is next.

OpenStudy (xapproachesinfinity):

my approach appears to be wrong! binomial distribution will work, yo most likely did \[P(X=5)=(1/4)^5(3/4)^4\] yes

OpenStudy (anonymous):

@xapproachesinfinity n=12, k=5 I used P(X=5) = 12C5 * \[0.25^{5}\] \[(1-0.25)^{12-5}\]=0.1032 is not 0.079

OpenStudy (anonymous):

@xapproachesinfinity 12C5* 0.25^5 (1-0.25)^12-5

OpenStudy (anonymous):

@perl We didn't learn geometric distribution. and I know how to use it either.

OpenStudy (perl):

Let F = fail to pick cherry candy S = successfully picks cherry candy X = how many times it takes to draw a candy until you get cherry P( X = 5) = P( pick cherry candy on fifth try) = P ( FFFFS) = P( F) * P(F) * P(F) * P(F) * P(S) = 9/12 * 9/12 * 9/12 * 9/12 * 3/12 =.0791

OpenStudy (anonymous):

@perl thanks a lot!!